consider the following three different investments. The cash flows generated by different alternatives are as follows:

End of Year A B C
1 300 0 600
2 300 0 0
3 300 0 0
4 300 600 600
5 300 600 0
6 300 600 600

Assuming a 15% annual discount rate, find the present value of each iinvestment.

(Show your work step by step)
Answer #1

A:

Present value = Annuity * [ 1 - 1 / ( 1 + r)n] / r

Present value = 300 * [ 1 - 1 / ( 1 + 0.15)6] / 0.15

Present value = 300 * 3.784483

Present value = $1,135.34

B:

Present value = 600 / ( 1 + 0.15)4 + 600 / ( 1 + 0.15)5 + 600 / ( 1 + 0.15)6

Present value = 343.0519 + 298.306 + 259.3966

Present value = $900.75

C:

Present value = 600 / ( 1 + 0.15)1 + 600 / ( 1 + 0.15)4 + 600 / ( 1 + 0.15)6

Present value = 521.7391 + 343.0519 + 259.3966

Present value = $1,124.19